# The inverse Kalman filter

### v5

@@ 1 Introduction
Dynamic linear models (DLMs) or linear state space models are ubiquitously used in modeling temporally correlated data [38, 5]. Each observation $y_{t}\in\mathbb{R}$ in DLMs is associated with a $q$-dimensional latent state vector $\bm{\theta}_{t}$, defined as:
@@
where $\mathbf{F}_{t}$ and $\mathbf{G}_{t}$ are matrices of dimensions $1\times q$ and $q\times q$, respectively, $\mathbf{W}_{t}$ is a $q\times q$ covariance matrix for $t=2,\dots,N$, and the initial state vector $\bm{\theta}_{1}\sim\mathcal{MN}(\mathbf{b}_{1},\mathbf{W}_{1})$ with $\mathbf{b}_{1}=\mathbf{0}$ assumed herein. The dependent structure of DLMs is illustrated in Fig. 1(a).
@@
Figure 1: The dependent structure of the dynamic linear model (panel (a)) and particle dynamics (panel (b)).
@@
To address this issue, we introduce the inverse Kalman filter (IKF), which computes $\mathbf{L}\mathbf{u}$ and $\mathbf{L}^{T}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$ operations without approximation, where $\mathbf{L}$ is the Cholesky factor of any $N\times N$ DLM-induced covariance $\bm{\Sigma}$ and $\mathbf{u}$ is an $N$-dimensional vector. The complexity is significantly smaller than $\mathcal{O}(N^{2})$ in direct computation. The dimension of the latent states $q$ does not exceed 3 in our applications, including commonly used models such as twice differentiable Gaussian processes with Matérn covariance in (15), often used as default in Gaussian process models.
@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in (1)-(2). This algorithm is applicable to any DLM specified in (1)-(2). Denote $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We develop fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th element of the vector of $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
For any $N$-dimensional vector ${\mathbf{u}}$, let $\tilde{x}_{N}={Q_{N}^{1/2}}u_{N}$, $\tilde{x}_{N-1}={Q_{N-1}^{1/2}}(\ell_{N,N-1}u_{N}+u_{N-1})$, and $\mathbf{g}_{N-1}=\mathbf{F}_{N}\mathbf{G}_{N}u_{N}$. For $t=N-2,\dots,1$, iteratively compute $\tilde{x}_{t}$ by
@@
where $\ell_{t+1,t}=\mathbf{F}_{t+1}\mathbf{G}_{t+1}\mathbf{K}_{t}$, ~$\tilde{\ell}_{t+1,t}={\mathbf{g}}_{t+1}\mathbf{G}_{t+1}\mathbf{K}_{t}$, and $\mathbf{K}_{t}$ is the Kalman gain
@@
with $\mathbf{B}_{t}=\mbox{cov}[\bm{\theta}_{t}\mid{\mathbf{y}}_{1:t-1}]$ and $Q_{t}=\mbox{var}[y_{t}\mid{\mathbf{y}}_{1:t-1}]$ from the Kalman filter in (A1) and (A2) in Appendix A, respectively. Then $\tilde{\mathbf{x}}=\mathbf{L}^{T}{\mathbf{u}}$.
@@
For any $N$-dimensional vector $\tilde{\mathbf{x}}$, let $x_{1}=\mathbf{F}_{1}{\mathbf{b}}_{1}+{Q_{1}}^{1/2}\tilde{x}_{1}$ and $\tilde{\mathbf{m}}_{1}={\mathbf{b}}_{1}+\mathbf{K}_{1}(x_{1}-\mathbf{F}_{1}{\mathbf{b}}_{1})$. For $t=2,\dots,N$, iteratively compute $x_{t}$ by
@@
where $Q_{t}$ and $\mathbf{K}_{t}$ are from (A2) and (6), respectively. Then $\mathbf{x}=\mathbf{L}\tilde{\mathbf{x}}$.
@@
The IKF algorithm, outlined in Algorithm 1, sequentially applies Lemmas 1 and 2 to reduce the computational cost and storage for $\bm{\Sigma}{\mathbf{u}}$ from $\mathcal{O}(N^{2})$ to $\mathcal{O}(q^{3}N)$ without approximation. The IKF applies to all DLMs, including Gaussian processes that admit equivalent DLM representation, such as Gaussian processes with Matérn kernels where roughness parameters $0.5$ and $2.5$ correspond to $q=1$ and $q=3$, respectively. Details are provided in the Supplementary Material.
Algorithm 1 Inverse Kalman filter (IKF) for computing $\bm{\Sigma}\mathbf{u}$.
1:$\mathbf{F}_{t},\mathbf{G}_{t},\mathbf{W}_{t},t=1,\dots,N$, an $N$-dimensional vector ${\mathbf{u}}$.
@@
Figure 2: Panels (a) and (b) compare the IKF (blue dots) with direct computation (red triangles) for computing $\bm{\Sigma}\mathbf{u}$ in the original time scale and logarithmic scale (base 10), respectively. Panel (c) shows the predictive error comparison between the non-robust IKF (light blue squares) and the robust IKF (blue dots).
While Algorithm 1 can be applied to DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in (1). This is because when $V_{t}=0$, $Q_{t}$ in (A2), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in (A3) due to unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the result of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
We compare the IKF approach with the direct matrix-vector multiplication of $\bm{\Sigma}\mathbf{u}$ in noise-free scenarios in Fig. 2. The experiments utilize $z_{t}=z(d_{t})$ with $d$ uniformly sampled from $[0,1]$ and the Matérn covariance with unit variance, roughness parameter $2.5$, and range parameter $\gamma=0.1$ [13]. Panels (a) and (b) show that the IKF significantly reduces computational costs compared to direct computation. The IKF achieves covariance matrix-vector multiplication for an output vector of $10^{6}$ dimensions in about 2 seconds on a desktop, making it highly scalable for iterative algorithms like the conjugate gradient algorithm [16] and the randomized log-determinant approximation [30]. Fig. 2(c) compares the maximum absolute error between the robust IKF with $V=0.1$ and the non-robust IKF with $V=0$. The non-robust IKF exhibits large numerical errors due to instability when $Q_{t}$ approaches zero, while the robust IKF remains stable by ensuring $Q_{t}$ is no smaller than $V$, even with near-singular covariance matrices $\bm{\Sigma}$.
Lemmas 1 and 2 facilitate the computation of each element in $\mathbf{L}$ using DLM parameters and enable the linear computation of $(\mathbf{L}^{T})^{-1}\tilde{\mathbf{x}}$. These results are summarized in Lemmas 3 and 4, respectively, with proofs provided in the Supplementary Material.
@@
where $\ell_{t^{\prime},t^{\prime}}=1$, and for $t^{\prime}>t$, $\ell_{t^{\prime},t}$ is defined as
@@
with $\prod^{t^{\prime}}_{l=t+1}\mathbf{G}_{l}=\mathbf{G}_{t^{\prime}}\mathbf{G}_{t^{\prime}-1}\dots\mathbf{G}_{t+1}$, $Q_{t}$ and $\mathbf{K}_{t}$ defined in (A2) and (6), respectively.
@@
For any $\tilde{\mathbf{x}}$, let $u_{N}={Q_{N}^{-1/2}}\tilde{x}_{N}$, and $u_{N-1}=Q_{N-1}^{-1/2}\tilde{x}_{N-1}-\ell_{N,N-1}u_{N}$. For $t=N-2,\dots,1$, iteratively compute $u_{t}$ by
@@
with ~${\tilde{\ell}}_{t+1,t}$ and $\ell_{t+1,t}$ defined in Lemma 1. Then ${\mathbf{u}}=(\mathbf{L}^{T})^{-1}\tilde{\mathbf{x}}$.
@@ 3 Nonparametric estimation of particle interaction functions
The IKF algorithm is motivated by the problem of estimating interactions between particles, which is crucial to understanding complex behaviors of molecules and active matter, such as migrating cells and flocking birds. Minimal physical models, such as the Vicsek model [36] and their extensions [1, 9], provide a framework for explaining the collective motion of active matter.
@@
where $\bm{\epsilon}_{i}\sim\mathcal{MN}(0,\sigma^{2}_{0}I_{D_{y}})$ denotes a Gaussian noise vector. The term $\mathbf{a}_{i,j,k}$ is a $D_{y}$-dimensional known factor loading that links the $i$th output to the $k$th neighbor in the $j$th interaction, with the unknown interaction function $z_{j}(\cdot)$ evaluated at a scalar-valued input $d_{i,j,k}$, such as the distance between particles $i^{\prime}$ and $k$, for $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, and $j=1,\dots,J$. For a dataset of $n_{p}$ particles over $n_{\tau}$ time points, the total number of observations is $\tilde{N}=nD_{y}=n_{p}n_{\tau}D_{y}$.
An illustrative example of this framework is an unnormalized Vicsek model. Unlike the original Vicsek model [36], which enforces a constant speed for all particles, we allow velocity magnitudes to vary over time. Time-varying velocities are more sensible in many real-world scenarios, including cellular movements studied in Section 4.4. When the system approaches a high degree of confluence, the average velocity of the cells typically decreases because of cell proliferation. Assume the 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, is updated by averaging the velocity of its neighbors from the previous time step. Specifically, for each direction $l=1,2$,
@@
where $\epsilon_{i^{\prime}}(\tau)$ is a zero-mean Gaussian noise with variance $\sigma^{2}_{0}$. The set of neighbors $ne_{i^{\prime}}(\tau-1)$ includes particles within a radius of $r$ from particle $i^{\prime}$ at time $\tau-1$, i.e., $ne_{i^{\prime}}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r\}$, where $\mathbf{s}_{i^{\prime}}(\tau-1)$ and $\mathbf{s}_{k}(\tau-1)$ are 2-dimensional position vectors of particle $i^{\prime}$ and its neighbor $k$, respectively, and $p_{i^{\prime}}(\tau-1)$ denotes the total number of neighbors of the $i^{\prime}$th particle, including itself, at time $\tau-1$. The illustration of this model is shown in Fig. S1(b) in the Supplementary Material.
@@
In this work, we model each latent interaction function $z_{j}(\cdot)$ nonparametrically using a Gaussian process. By utilizing an equivalent DLM representation, we can leverage the proposed IKF algorithm to significantly expedite computation. This includes commonly used kernels such as Matérn kernel with a half-integer roughness parameter [13], often used as the default setting for Gaussian process models to predict nonlinear functions, as the smoothness of the process can be controlled by its roughness parameters [12].
For each interaction $j$, we form the input vector $\mathbf{d}^{(u)}_{j}=[d^{(u)}_{1,j},\dots,d^{(u)}_{N_{j},j}]^{T}$, where ‘u’ indicates that the input entries are unordered, $d^{(u)}_{t,j}=d_{i,j,k}$ with $t=\sum^{i-1}_{i^{\prime}=1}p_{i^{\prime},j}+k$ for any tuple $(i,j,k)$, $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, $j=1,\dots,J$, and $N_{j}=\sum^{n}_{i=1}p_{i,j}$. The marginal distribution of the $j$th factor follows $\mathbf{z}_{j}^{(u)}=[z_{j}(d_{1,j}^{(u)}),\dots,z_{j}(d_{N_{j},j}^{(u)})]^{T}\sim\mathcal{MN}(\mathbf{0},\,\bm{\Sigma}_{j}^{(u)})$, where the $(t,t^{\prime})$th entry of the $N_{j}\times N_{j}$ covariance matrix $\bm{\Sigma}_{j}^{(u)}$ is $\mbox{cov}[z_{j}(d_{t,j}^{(u)}),z_{j}(d_{t^{\prime},j}^{(u)})]=\sigma^{2}_{j}c_{j}(d_{t,j}^{(u)},d_{t^{\prime},j}^{(u)})=\sigma^{2}_{j}c_{j}(d)$, with $d=|d_{t,j}^{(u)}-d_{t^{\prime},j}^{(u)}|$, $\sigma^{2}_{j}$ and $c_{j}(\cdot)$ being the variance parameter and the correlation function, respectively. We employ the Matérn correlation with roughness parameters $\nu_{j}=0.5$ and $\nu_{j}=2.5$. For $\nu_{j}=0.5$, the Matérn correlation is the exponential correlation $c_{j}(d)=\exp(-d/\gamma_{j})$, while for $\nu_{j}=2.5$, the Matérn correlation follows:
@@
$\left(\mathbf{y}\mid\bm{\Sigma}^{(u)}_{j},\sigma^{2}_{0}\right)\sim\mathcal{MN}\left(\mathbf{0},\sum^{J}_{j=1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{A}^{T}_{j}+\sigma^{2}_{0}\mathbf{I}_{\tilde{N}}\right),$ (16)
where $\mathbf{A}_{j}$ is a sparse $\tilde{N}\times N_{j}$ block diagonal matrix, such that the $i$th diagonal block is a $D_{y}\times p_{i,j}$ matrix $\mathbf{A}_{i,j}=[\mathbf{a}_{i,j,1},\dots,\mathbf{a}_{i,j,p_{i,j}}]$ for $i=1,\dots,n$. The posterior predictive distribution of the latent variable $z_{j}(d^{*})$ for any given input $d^{*}$ is also a normal distribution:
@@
where the predictive mean and variance are given by

### v6

@@ 1 Introduction
Dynamic linear models (DLMs) or linear state space models are ubiquitously used in modeling temporally correlated data [38, 5]. Each observation $y_{t}\in\mathbb{R}$ in DLMs is associated with a latent state vector $\bm{\theta}_{t}\in\mathbb{R}^{q}$, defined as
@@
where $\mathbf{F}_{t}$ and $\mathbf{G}_{t}$ are matrices of dimensions $1\times q$ and $q\times q$, respectively, $\mathbf{W}_{t}$ is a $q\times q$ covariance matrix for $t=2,\dots,N$ and the initial state vector follows a multivariate normal distribution $\bm{\theta}_{1}\sim\mathcal{MN}(\mathbf{b}_{1},\mathbf{W}_{1})$ with $\mathbf{b}_{1}=\mathbf{0}$ assumed herein. The dependent structure of the DLM is illustrated in Fig. 1(a).
@@
Figure 1: The dependent structure of the (a) dynamic linear model and (b) particle dynamics.
@@
To address this issue, we introduce the inverse Kalman filter (IKF), which computes $\mathbf{L}\mathbf{u}$ and $\mathbf{L}^{T}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$ operations without approximation, where $\mathbf{L}$ is the Cholesky factor of any $N\times N$ DLM-induced covariance $\bm{\Sigma}$ and $\mathbf{u}$ is an $N$-dimensional vector. The complexity is significantly smaller than $\mathcal{O}(N^{2})$ in direct computation. The dimension of the latent states $q$ does not exceed 3 in our applications, and this choice includes many commonly used models such as twice differentiable Gaussian processes with Matérn covariance in (15), often assumed as default in Gaussian process models.
@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in (1)-(2). This algorithm is applicable to any DLM specified in (1)-(2). Define $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We develop fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2 below, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th elements of the vectors $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
For any $N$-dimensional vector $\mathbf{u}$}, let $\tilde{x}_{N}={Q_{N}^{1/2}}u_{N}$, $\tilde{x}_{N-1}={Q_{N-1}^{1/2}}(\ell_{N,N-1}u_{N}+u_{N-1})$, and $\mathbf{g}_{N-1}=\mathbf{F}_{N}\mathbf{G}_{N}u_{N}$. For $t=N-2,\dots,1$, iteratively compute $\tilde{x}_{t}$ using
@@
with $\ell_{t+1,t}=\mathbf{F}_{t+1}\mathbf{G}_{t+1}\mathbf{K}_{t}$, $\tilde{\ell}_{t+1,t}={\mathbf{g}}_{t+1}\mathbf{G}_{t+1}\mathbf{K}_{t}$, and the Kalman gain
@@
where $\mathbf{B}_{t}=\mbox{cov}[\bm{\theta}_{t}\mid{\mathbf{y}}_{1:t-1}]$ and $Q_{t}=\mbox{var}[y_{t}\mid{\mathbf{y}}_{1:t-1}]$ are respectively defined in the Kalman filter in (A1) and (A2) in Appendix A. Then $\tilde{\mathbf{x}}=\mathbf{L}^{T}{\mathbf{u}}$.
@@
For any $N$-dimensional vector $\tilde{\mathbf{x}}$, let $x_{1}=\mathbf{F}_{1}{\mathbf{b}}_{1}+{Q_{1}}^{1/2}\tilde{x}_{1}$ and $\tilde{\mathbf{m}}_{1}={\mathbf{b}}_{1}+\mathbf{K}_{1}(x_{1}-\mathbf{F}_{1}{\mathbf{b}}_{1})$. For $t=2,\dots,N$, iteratively compute $x_{t}$ using
@@
where $\mathbf{K}_{t}$ and $Q_{t}$ are respectively defined in (6) and (A2) in the Appendix. Then $\mathbf{x}=\mathbf{L}\tilde{\mathbf{x}}$.
@@
The IKF algorithm, outlined in Algorithm 1, sequentially applies Lemmas 1 and 2 to reduce the computational cost and storage for $\bm{\Sigma}{\mathbf{u}}$ from $\mathcal{O}(N^{2})$ to $\mathcal{O}(q^{3}N)$ without approximation. Note that the $\mathcal{O}(q^{3}N)$ is primarily due to the matrix-matrix multiplication required to obtain $\mathbf{B}_{t}$ (Eq. (A1)) in the initial KF. Once these quantities are pre-computed, the core iterations in Lemmas 1-2 only involve matrix-vector multiplications, which require $\mathcal{O}(q^{2}N)$. The IKF applies to all DLMs, including Gaussian processes that admit equivalent DLM representation, such as Gaussian processes with Matérn kernels where roughness parameters $0.5$ and $2.5$ correspond to $q=1$ and $q=3$, respectively. Details are provided in the Supplementary Material.
Algorithm 1 The IKF for computing $\bm{\Sigma}\mathbf{u}$
[…]
@@
Figure 2: Panels (a) and (b) compare the IKF (blue dots) with direct computation (red triangles) for calculating $\bm{\Sigma}\mathbf{u}$ in the original time scale and logarithmic scale (base 10), respectively. Panel (c) shows the predictive error comparison between the non-robust IKF (light blue squares) and the robust IKF (blue dots).
While Algorithm 1 can be applied to a DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in (1). This is because, when $V_{t}=0$, $Q_{t}$ in (A2), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in (A3) due to the unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the result of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in the KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
We compare the IKF approach with the direct matrix-vector multiplication of $\bm{\Sigma}\mathbf{u}$ in noise-free scenarios in Fig. 2. The inputs of $\bm{\Sigma}$ are uniformly sampled from $[0,1]$ and the Matérn covariance with unit variance, roughness parameter $2.5$, and range parameter $\gamma=0.1$ [13] is used. Panels (a) and (b) show that the IKF significantly reduces computational costs compared to direct computation. The IKF achieves covariance matrix-vector multiplication for an output vector of $10^{6}$ dimensions in about 2 seconds on a desktop, making it highly scalable to be applied in iterative algorithms like the conjugate gradient algorithm [16] and the randomized log-determinant approximation [30]. Fig. 2(c) compares the maximum absolute error between the robust IKF with $V=0.1$ and the non-robust IKF with $V=0$. The non-robust IKF exhibits large numerical errors due to instability when $Q_{t}$ approaches zero, while the robust IKF remains stable by ensuring that $Q_{t}$ is no smaller than $V$, even with near-singular covariance matrices $\bm{\Sigma}$.
Lemmas 1 and 2 facilitate the computation of each element in $\mathbf{L}$ using DLM parameters and enable the linear computation of $(\mathbf{L}^{T})^{-1}\tilde{\mathbf{x}}$. These results are respectively summarized in Lemmas 3 and 4 below, with proofs provided in the Supplementary Material.
@@
[…]
@@
with $\prod^{t^{\prime}}_{l=t+1}\mathbf{G}_{l}=\mathbf{G}_{t^{\prime}}\mathbf{G}_{t^{\prime}-1}\dots\mathbf{G}_{t+1}$, $\mathbf{K}_{t}$ and $Q_{t}$ respectively defined in (6) and (A2).
@@
For any $\tilde{\mathbf{x}}$, let $u_{N}={Q_{N}^{-1/2}}\tilde{x}_{N}$, and $u_{N-1}=Q_{N-1}^{-1/2}\tilde{x}_{N-1}-\ell_{N,N-1}u_{N}$. For $t=N-2,\dots,1$, iteratively compute $u_{t}$ using
@@
with ${\tilde{\ell}}_{t+1,t}$ and $\ell_{t+1,t}$ defined in Lemma 1. Then ${\mathbf{u}}=(\mathbf{L}^{T})^{-1}\tilde{\mathbf{x}}$.
@@ 3 Nonparametric estimation of particle interaction functions
The IKF algorithm is motivated by the problem of estimating interactions between particles, which is crucial to understanding complex behaviors of molecules and active matter, such as migrating cells and flocking birds. Minimal physical models, such as the Vicsek model [36] and its extensions [1, 9], provide a framework for explaining the collective motion of active matter.
@@
with $\bm{\epsilon}_{i}\sim\mathcal{MN}(0,\sigma^{2}_{0}I_{D_{y}})$ denoting a Gaussian noise vector. The term $\mathbf{a}_{i,j,k}$ is a $D_{y}$-dimensional known factor loading that links the $i$th output to the $k$th neighbor in the $j$th interaction, with the unknown interaction function $z_{j}(\cdot)$ evaluated at a scalar-valued input $d_{i,j,k}$, such as the distance between particles $i^{\prime}$ and $k$, for $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, and $j=1,\dots,J$. For a dataset of $n_{p}$ particles over $n_{\tau}$ time points, the total number of observations is $\tilde{N}=nD_{y}=n_{p}n_{\tau}D_{y}$.
An illustrative example of this framework is an unnormalized Vicsek model. Unlike the original Vicsek model [36], which enforces a constant speed for all particles, we allow velocity magnitudes to vary over time. Time-varying velocities are more sensible in many real-world scenarios, including cellular movements studied in Section 4.4. When the system approaches a high degree of confluence, the average velocity of the cells typically decreases because of cell proliferation. Assume that the 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, is updated by averaging the velocity of its neighbors from the previous time step. Specifically, for each direction $l=1,2$,
@@
[…]
@@
In this work, we model each latent interaction function $z_{j}(\cdot)$ nonparametrically using a Gaussian process. By utilizing an equivalent DLM representation, we can leverage the proposed IKF algorithm to significantly expedite computation. This includes commonly used kernels such as the Matérn kernel with a half-integer roughness parameter [13], often used as the default setting for Gaussian process models to predict nonlinear functions, as the smoothness of the process can be controlled by its roughness parameters [12].
For each interaction $j$, we form the input vector $\mathbf{d}^{(u)}_{j}=[d^{(u)}_{1,j},\dots,d^{(u)}_{N_{j},j}]^{T}$, where superscript $(u)$ indicates that the input entries are unordered, $d^{(u)}_{t,j}=d_{i,j,k}$ with $t=\sum^{i-1}_{i^{\prime}=1}p_{i^{\prime},j}+k$ for any tuple $(i,j,k)$, $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, $j=1,\dots,J$, and $N_{j}=\sum^{n}_{i=1}p_{i,j}$. The marginal distribution of the $j$th factor follows $\mathbf{z}_{j}^{(u)}=[z_{j}(d_{1,j}^{(u)}),\dots,z_{j}(d_{N_{j},j}^{(u)})]^{T}\sim\mathcal{MN}(\mathbf{0},\,\bm{\Sigma}_{j}^{(u)})$, where the $(t,t^{\prime})$th entry of the $N_{j}\times N_{j}$ covariance matrix $\bm{\Sigma}_{j}^{(u)}$ is $\mbox{cov}[z_{j}(d_{t,j}^{(u)}),z_{j}(d_{t^{\prime},j}^{(u)})]=\sigma^{2}_{j}c_{j}(d_{t,j}^{(u)},d_{t^{\prime},j}^{(u)})=\sigma^{2}_{j}c_{j}(d)$, with $d=|d_{t,j}^{(u)}-d_{t^{\prime},j}^{(u)}|$, and $\sigma^{2}_{j}$ and $c_{j}(\cdot)$ the variance parameter and correlation function, respectively. We employ the Matérn correlation with roughness parameters $\nu_{j}=0.5$ and $\nu_{j}=2.5$. For $\nu_{j}=0.5$, the Matérn correlation is the exponential correlation $c_{j}(d)=\exp(-d/\gamma_{j})$, while for $\nu_{j}=2.5$, the Matérn correlation follows
@@
By integrating out latent factor processes $\mathbf{z}_{j}$, the marginal distribution of the observational vector $\mathbf{y}=(\mathbf{y}^{T}_{1},\dots,\mathbf{y}^{T}_{n})^{T}$, with dimension $\tilde{N}=nD_{y}$, follows a multivariate normal distribution:
@@
Here, $\mathbf{A}_{j}$ is a sparse $\tilde{N}\times N_{j}$ block diagonal matrix, such that the $i$th diagonal block is a $D_{y}\times p_{i,j}$ matrix $\mathbf{A}_{i,j}=[\mathbf{a}_{i,j,1},\dots,\mathbf{a}_{i,j,p_{i,j}}]$ for $i=1,\dots,n$. The posterior predictive distribution of the latent variable $z_{j}(d^{*})$ for any given input $d^{*}$ is also a normal distribution:
@@
with the predictive mean and variance given by